OpenStudy (anonymous):

cscxcotx(1-cos^2x)=___x

OpenStudy (anonymous):

\[ \csc x\cot x(1-\cos^2x)= \frac 1{\sin x} \frac {\cos x}{\sin x} \sin^2 x=\cos x \]